Problem 5. (8 points) Consider rolling a six-sided die. Let A be the set of outcomes where the roll is an even number Let B be the set of outcomes where the roll is greater than 3. Calculate the sets on both sides of De Morgans laws and verify that the equality holds. (AUB)c A n B
Answer #1

Using definitions of A and B we have

A={2,4,6} , B={4,5,6} and А 1.3.5 . Вс 1.2.3

So, A U B 12, 4, 5, 6} ​​​​​​, and A^ccap B^c={1,3}~,~(Acup B) ^c={1,3} ​​​​​​

Therefore, (Acup B) ^c=A^ccap B^c ​​​​​​

Also, Acap B={4,6} ​​​​​​, and  (AnBF= {1.2, 3.5} Ac U B-(1.2.3.5} ,

Hence ​(Acap B) ^c=A^ccup B^c